The product of a mans age 6 years ago and his age in 10 years is 960. Find his age now
Answer by Cromlix(4381)

a mans age 6 years ago
(x - 6)
and his age in 10 years
(x + 10)
The product of a mans age 6 years ago and his age in 10 years is 960
(x - 6)(x + 10) = 960
x^2 + 4x - 60 = 960
x^2 + 4x - 1020 = 0
(x - 30)(x + 34) = 0
x + 34 = 0
x = -34 (no answer as x < 0)
x - 30 = 0
x = 30.
The man's age is 30 years old.
Hope this helps.
:-)